A 100-watt motor (60% efficient) is available to r
Answer #1

Efficiency

n=\frac{P_{out}}{P_{in}}

if

the power supplied to the motor

P_{in}=100W
efficiency

n=0.60

finally

n=\frac{P_{out}}{P_{in}}\Rightarrow P_{out}=P_{in}\cdot n=100W\cdot 0.60=60W

the potency is equal to

P_{out}=\frac{E}{t}

where:

energy is equal to the potential energy

E=m\cdot g\cdot h
time

t=65s
mass expression

P_{out}=\frac{m\cdot g\cdot h}{t}\Rightarrow m=\frac{P_{out}\cdot t}{ g\cdot h}

evaluating numerically

m=\frac{60W\cdot 65s}{ 9.81\frac{m}{s^{2}}\cdot 5m}=79.5kg